OpenStudy (anonymous):

0 = x2 + 10x + 5 how to solve this equation by completing the square and using the quadratic formula?

OpenStudy (anonymous):

\[x^2+10x=-5\] \[(x+5)^2=-5+25=20\] \[x+5=\pm\sqrt{20}=\pm2\sqrt{5}\] \[x=-5\pm2\sqrt{5}\]